OpenStudy (anonymous):

Let f(x) = x2 – 7x + 3 and g(x) = x + 4. Find f(x) + g(x). x^2 - 8x + 7 x^2 - 6x + 7 x^3 - 3x + 7 x^3 - 7x + 7

OpenStudy (anonymous):

x2 – 7x + 3+ x + 4

OpenStudy (anonymous):

x^2-6*x+7

OpenStudy (anonymous):

f(x) + g(x) = (x2 - 7x + 3) + (x+4) = x2 -7x +3 +x +4 =x2 -6x +7

OpenStudy (anonymous):

Janie29 is correct